OpenStudy (anonymous):

Slove the system by substitution. 3x+y=2 4x+y=20

OpenStudy (anonymous):

3x+y=2 y=2-3x (subtracted 3x from both sides to get Y by it self) 4x+(2-3x)=20 (substituted in for Y) 4x-3x=18 (subtracted the 2) x=18 ( 4x-3x gives u x) 4(18)+y=20 (plug in what u got for X an then solve for Y) 72+y=20 (multiplied 4 x 18 and got 72) y=-52 ( subtracted 20-72=-52)